# The Manakov-Zakharov-Ward model as an integrable decoupling limit of the membrane

### v1

@@ 1 Introduction
A non-relativistic decoupling limit, as proposed here, fits into a recently observed pattern that certain decoupling limits that isolate BPS $p$‑brane dynamics and associated matrix models [9, 10, 11], naturally giving rise to non‑Lorentzian target‑space geometries known as $p$-brane Newton-Cartan geometry (see [12] for a recent review). More recently, this has also been discussed extended to eleven-dimensional supergravity [13, 14, 15, 16].
@@ 1.1 The bosonic membrane
$S$ $\sim\int\mathrm{d}^{3}\sigma\sqrt{\gamma}\left(\frac{1}{2}\gamma^{MN}\partial_{M}X^{\mu}\partial_{N}X^{\nu}g_{\mu\nu}\right.$ (1.1)
${}\qquad+\left.\frac{1}{3!}\epsilon^{KLM}\partial_{K}X^{\mu}\partial_{L}X^{\nu}\partial_{M}X^{\rho}C_{\mu\nu\rho}+\frac{1}{2}\right)$
in a bosonic supergravity background described by a metric $g$ and three-form gauge field $C$. $\sigma^{M}$ are the coordinates on the membrane world-volume, $X^{\mu}(\sigma)$ the embedding coordinate fields, and $\gamma_{MN}$ is the (dynamical) world-volume metric. The cosmological constant term, as usual, ensures that the world-volume metric becomes the induced metric (as algebraic equation of motion for the world-volume metric $\gamma$):
$\gamma_{MN}=\partial_{M}X^{\mu}\partial_{N}X^{\nu}g_{\mu\nu}.$ (1.2)
In contrast to the string, reparameterisation invariance is not sufficient to gauge-fix it to be the flat metric. The equations of motion for the world-volume coordinate fields $X^{\mu}(\sigma)$ are
00 0 $=\partial_{M}(\gamma^{MN}\partial_{N}X^{\mu})+\gamma^{MN}\Gamma^{\mu}{}_{\kappa\lambda}\partial_{M}X^{\kappa}\partial_{N}X^{\lambda},$ (1.3)
${}\quad+\epsilon^{NKL}\frac{1}{3!}{F^{\mu}}_{\nu\kappa\lambda}\partial_{N}X^{\nu}\partial_{K}X^{\kappa}\partial_{L}X^{\lambda},$
@@
Famously [17, 18], even this supposedly most simple case leads to complications, such as a continuous spectrum and spike instabilities. Quantum non-integrability of the membrane in the sense of scattering of fluctuations around a classical solution, and its relation to string theory, has recently been shown in [19]. In this letter, we will see that a large wrapping limit together with a scaling limit of the background (or alternatively: a non-relativistic limit of fluctuations around a wrapped membrane) can simplify these equations. In the special case considered here, the result will even be a classically integrable system.
@@ 1.2 The Manakov-Zakharov-Ward model
Classically integrable field theories in dimensions higher than $(1+1)$ are rare, especially those of the $\sigma$-model type, which are abundant in $(1+1)$ dimensions (see [20, 21] for recent reviews). A notable exception is the Manakov-Zakharov-Ward (MZW) model [7, 8], an integrable field theory in $(1+2)$ dimensions. It arises via dimensional reduction from the anti-self-dual Yang-Mills equations in $(2+2)$-dimensional spacetime. This reduction involves assuming that the gauge potentials are independent of one spatial coordinate, resulting in a non-Lorentz invariant system in $(1+2)$ dimensions.
@@
1.3 A non-relativistic expansion of eleven-dimensional supergravity
The non-relativistic nature of the MZW-model suggests the use of non-relativistic geometry. In particular, we are interested in the application to membranes in eleven-dimensional supergravity. A finite non-relativistic limit of eleven-dimensional supergravity has been obtained in [30, 31] and gives rise to the following expansions of metric $g$ and three-form gauge field $C$ as series of some dimensionless parameter $c\rightarrow\infty$:
$\hat{g}_{\mu\nu}$ $=c^{2}\tau_{\mu}{}^{A}\tau_{\nu}{}^{B}\eta_{AB}+c^{-1}H_{\mu\nu}+\mathcal{O}(c^{-4}),$ (1.5)
$\hat{C}_{\mu_{1}\mu_{2}\mu_{3}}$ $=c^{3}\tau_{\mu_{1}}{}^{A}\tau_{\mu_{2}}{}^{B}\tau_{\mu_{3}}{}^{C}\epsilon_{ABC}+C_{\mu_{1}\mu_{2}\mu_{3}}+\mathcal{O}(c^{-3}).$ (1.6)
$\eta_{AB}$ and $\epsilon_{ABC}$ are the $(1+2)$-dimensional Minkowski metric and totally antisymmetric tensor. The geometry, locally, describes a split into a $(1+2)$-dimensional relativistic and an eight-dimensional non-relativistic space, by three longitudinal (so-called clock) one-formd $\tau^{A}$ and a transverse metric $H$. In inverse of these objects is defined by the completeness and orthogonality conditions
$H^{\mu\rho}H_{\rho\nu}+{\tau^{\mu}}_{A}\tau_{\nu}{}^{A}$ $=\delta_{\nu}^{\mu},$
The relevant examples in this letter are very simplistic, allowing a global split into longitudinal and transverse space with coordinates $x^{\mu}=(x^{M},x^{m})$. Moreover, $\mathrm{d}\tau^{A}=0$, the condition for the so-called Newton-Cartan torsionless membrane geometry, will always be satisfied here.
Similarly to the non-relativistic expansion of the bosonic string [32, 33], a consistent treatment of the membrane in a Newton-Cartan background requires the expansion of the coordinate fields. Here, we assume
$\hat{X}^{\mu}(\sigma)=Y^{\mu}(\sigma)+c^{-3}X^{\mu}(\sigma)+\mathcal{O}(c^{-6}).$ (1.7)
Although this expansion has not been discussed in the literature, [^1] There is only sparse literature on the non-relativistic membrane world-volume. In [34] an $c^{-2}$-expansion for the world-volume fields has been assumed in addition to the $c^{-3}$-expansion of the target space geometry, essentially leading to a resulting expansion at all orders in $c^{-1}$. it coincides with the $c^{-3}$-expansion of the target space and will be crucial in what is to follow.
2 Two decoupling limits for the membrane on T2
@@
Here, we present two (ultimately equivalent) views on a decoupling limit that can be interpreted either as a large wrapping limit around the $T^{2}$ or a non-relativistic expansion. In particular, we aim at the following, at leading order in these limits:
@@
  • membrane dynamics along $\mathcal{M}_{H}$rightarrow MZW-model
@@
2.1 Post-Newtonian dynamics
Let us propose the following expansion for the background geometry to (2)
$\hat{\mathrm{d}s}^{2}$ $=c^{2}\eta_{MN}\mathrm{d}x^{M}\mathrm{d}x^{N}+c^{-1}H_{mn}\mathrm{d}x^{m}\mathrm{d}x^{n}$ (2.1)
$\hat{F}$ $=F=\frac{1}{3!}F_{klmN}\mathrm{d}x^{k}\wedge\mathrm{d}x^{l}\wedge\mathrm{d}x^{m}\wedge\mathrm{d}x^{N}$ (2.2)
where coordinate $x^{\mu}=(x^{M},x^{m})$ are chosen that make the split into $\mathcal{M}_{\tau}$ and $\mathcal{M}_{H}$ manifest. [^3] Indices of tensor of expansions (without $\hat{}$ are raised and lowered by $\eta$ or $H$. Compared to (1.5), this corresponds to the choice of clock forms $\tau^{M}=\mathrm{d}x^{M}$, which in turn causes the $c^{3}$-contribution to $\hat{F}=\mathrm{d}\hat{C}$ from (1.6) to vanish.
The striven for decoupling of three-dimensional gravity on the membrane world-volume is achieved at non-trivial leading orders of the following expansion for the coordinate fields $\hat{X}^{\mu}(\sigma)$
$\hat{X}^{\mu}(\sigma)$ $=(\hat{X}^{M}(\sigma),\hat{X}^{m}(\sigma))$
$=\left(\sigma^{M}+c^{-3}X^{M}(\sigma),c^{-3}X^{m}(\sigma)\right).$ (2.3)
We assume that $X^{M}(\sigma)$ is periodic in $\sigma^{M}$ for $M=1,2$. The fields $X^{\mu}=(X^{M},X^{m})$ correspond to fluctuations around a static membrane in $\mathbb{R}_{t}\times T^{2}$ that is wrapped around the $T^{2}$. These fluctuations are suppressed by the scale $\epsilon\sim\frac{1}{c^{3}}$. As desired, this expansion decouples the world-volume gravity at leading orders in $\frac{1}{c^{3}}$. The inverse of induced metric $\hat{\gamma}_{MN}=\partial_{M}\hat{X}^{\mu}\partial_{N}\hat{X}^{\nu}\hat{g}_{MN}$ then has the following expansion
$\hat{\gamma}^{MN}=c^{-2}\eta^{MN}-c^{-5}2\partial^{(M}X^{N)}+\mathcal{O}(c^{-8})$ (2.4)
At leading orders indices $M,N,...$ are raised with $\eta$. Moreover, this expansion ensures that the membrane couples (at leading order) only to the leading-order background fields. Moreover, the resulting theory in the decoupling limit is a three-dimensional non-relativistic (non-Lorentz invariant) theory, as is striven for in order to reproduce the MZW-model. The equations motion at leading order are:
00 0 $=\frac{1}{c^{5}}\left(-\partial^{M}\partial_{N}X^{N}+\mathcal{O}\left(c^{-3}\right)\right.$
${}\quad\left.+\frac{1}{c^{6}}\frac{1}{3!}\eta^{MN}F_{Nklm}\partial_{K}X^{k}\partial_{L}X^{l}\partial_{R}X^{m}\epsilon^{KLR}\right)$ (2.5)
00 0 $=\frac{1}{c^{5}}\Big{(}\partial_{M}\partial^{M}X^{m}+...$
${}\quad\left.+\frac{1}{2}H^{mn}F_{klnN}\partial_{K}X^{k}\partial_{L}X^{l}\epsilon^{KLN}+\mathcal{O}\left(c^{-3}\right)\right)$ (2.6)
where ’...’ denotes the Christoffel symbol contribution that will not be relevant in the following. Hence, the four-form flux that mixes longitudinal and transversal degrees of freedom decouples from the longitudinal $X^{M}$ equations of motion.
The leading part of the longitudinal coordinates’ $X^{M}$ dynamics comes from a non-trivial combination the next-to-leading order of the membrane coordinate (2.3) and world-volume metric (2.4) expansions. The resulting powers of $\frac{1}{c}$ in front of the four-form flux are combinations of $\frac{1}{c^{3}}$ factors of the fluctuations $X^{\mu}$ in (2.3) and the effect of raising the index of $F$ with the metric (2.1).
2.2 A large wrapping limit
An alternative point of view can be provided without resorting to membrane Newton-Cartan geometry and non-relativistic interpretation by considering the more general ansatz introducing scales $R$ and $\bar{R}$ for the longitudinal and transversal spaces: [^4] The scaling like this for $F$ has to be introduced in order to be consistent with the supergravity equations. The choice taken here is not general, but the one that will appear below.
$\hat{\mathrm{d}s}^{2}$ $=R^{2}\eta_{MN}\mathrm{d}x^{M}\mathrm{d}x^{N}+\bar{R}^{2}H_{mn}\mathrm{d}x^{m}\mathrm{d}x^{n},$ (2.7)
$\hat{F}$ $=F=\frac{1}{3!}R^{2}\bar{R}F_{klmN}\mathrm{d}x^{k}\wedge\mathrm{d}x^{l}\wedge\mathrm{d}x^{m}\wedge\mathrm{d}x^{N}.$ (2.8)
In contrast to the non-relativistic setting, we look at arbitrary fluctuations around a static wrapped membrane in the large wrapping number limit $\omega\rightarrow\infty$:
@@ 2.2 A large wrapping limit
The parameter $\omega$ describes both a ’world-volume speed of light’ and the wrapping number around the torus. [^5] In principle, these parameters could be different, of course – i.e. $X^{M}=(\omega_{0}\sigma^{0},\omega_{1}\sigma^{1},\omega_{2}\sigma^{2})$. For the limit in question, we want to assume $\omega_{0}\sim\omega_{1}\sim\omega_{2}\sim\omega\rightarrow\infty$. Finite quotients could be compensated for by rescaling of the membrane coordinates $\sigma^{M}$. In the limit $\omega\rightarrow\infty$, the $\mathcal{O}(\omega^{0})$ world-volume fields $x^{\mu}(\sigma)$ are suppressed and the induced metric simplifies significantly. In particular, one can compute the inverse
$\gamma^{MN}=\frac{1}{\omega^{2}R^{2}}\left(\eta^{MN}-\frac{1}{\omega}2\partial^{(M}x^{N)}+\mathcal{O}(\omega^{-2})\right).$ (2.10)
in order to compute the expansion of the equations of motion in leading order
00 0 $=-\frac{1}{\omega^{2}R^{2}}\left(\partial^{K}\partial_{M}X^{M}+\mathcal{O}\left(\frac{1}{\omega}\right)\right)$ (2.11)
${}\quad+\frac{\bar{R}^{2}}{R}\frac{1}{3!}\eta^{KM}F_{Mklm}\partial_{K}X^{k}\partial_{L}X^{l}\partial_{R}X^{m}\epsilon^{KLR},$
00 0 $=\frac{1}{\omega^{2}R^{2}}\left(\partial_{M}\partial^{M}X^{m}+...\right)$ (2.12)
${}\quad+\frac{1}{2}R\eta^{mn}F_{klnN}\partial_{K}X^{k}\partial_{L}X^{l}\epsilon^{KLN}+\mathcal{O}\left(\frac{1}{\omega^{3}R^{2}}\right).$
Requiring that the four-form flux couples to the leading order transverse dynamics and decoupling of the longitudinal and transverse degrees of freedom at leading order leads to the following scaling for the (dimensionless) scales $R$ and $\bar{R}$:
$R^{3}\sim\frac{1}{\omega^{2}}\rightarrow 0,\qquad\mathcal{O}(\bar{R}^{2})<\mathcal{O}\left(\frac{1}{\omega^{2}R}\right).$ (2.13)
The results of the two decoupling limits – the post-Newtonian and the large wrapping one – agree for $\omega\sim c^{3}$, $R\sim c$, and $\bar{R}\sim\frac{1}{\sqrt{c}}$.
@@ 3.1 A toy model background
In order to obtain the MZW-model from the transverse equations of motion at leading order, like (2.6), we specify $\mathcal{M}_{H}$ to be a Lie group $G$ [^6] For now, we are not concerned with dimensionality of the background. The background expansion in the conventions for the non-relativstic limit are
$\hat{\mathrm{d}s}^{2}$ $=c^{2}\eta_{MN}\mathrm{d}x^{M}\mathrm{d}x^{N}+c^{-1}\kappa_{ab}e^{a}e^{b}$ (3.1)
$\hat{F}$ $=F=-\frac{1}{3!}f_{abc}e^{a}\wedge e^{b}\wedge e^{c}\wedge v.$
$e^{a}={e_{m}}^{a}\mathrm{d}x^{m}$ are the Maurer-Cartan one-forms on $G$, $f_{abc}$ are the structure constants for the Lie algebra $\mathfrak{g}$ of $G$, and $\kappa$ is the Killing form on $\mathfrak{g}$. $v$ is a one form on $\mathcal{M}_{\tau}$. We choose $v=\mathrm{d}x^{0}$ or $v=\mathrm{d}x^{1/2}$. This choice will be related to the choice of $v$ in the MZW-model (1.4). In the same way as in the previous section, at leading order the membrane equations of motion, now expressed in terms of currents $j^{a}={e_{m}}^{a}\mathrm{d}X^{m}$, are
00 0 $=\partial_{M}\partial_{N}X^{N},$ (3.2)
00 0 $=\partial^{N}j_{N}^{a}-\frac{1}{2}{f^{a}}_{bc}j^{b}_{M}j^{c}_{N}v_{L}\epsilon^{MNL}.´,$ (3.3)
for the expansion (2.3). We recognise the MZW equation of motion (1.4) for the transverse directions. The longitudinal equations of motion $\partial_{M}\partial_{N}X^{N}=0$ decouple, but are solvable in their own right. For the purpose of this paper, we will assume the solution $X^{N}=0$ in the following.

### v2

@@ 1 Introduction
A non-Lorentzian decoupling limit, as proposed here, fits into a recently observed pattern that certain decoupling limits that isolate BPS pp‑brane dynamics and associated matrix models [9, 10, 11], naturally giving rise to non‑Lorentzian target‑space geometries known as $p$-brane Newton-Cartan geometry (see [12] for a recent review). More recently, this has also been discussed extended to eleven-dimensional supergravity [13, 14, 15, 16].
@@ 1.1 The bosonic membrane
$S$ $\sim\int\mathrm{d}^{3}\sigma\left(\frac{1}{2}\sqrt{-\gamma}\gamma^{MN}\partial_{M}\hat{X}^{\mu}\partial_{N}\hat{X}^{\nu}g_{\mu\nu}\right$. (1.1)
${}\qquad+\left.\frac{1}{3!}\epsilon^{KLM}\partial_{K}\hat{X}^{\mu}\partial_{L}\hat{X}^{\nu}\partial_{M}\hat{X}^{\rho}C_{\mu\nu\rho}-\frac{1}{2}\sqrt{-\gamma}\right)$
in a bosonic supergravity background described by a metric $g$ and three-form gauge field $C$. $\sigma^{M}$ are the coordinates on the membrane world-volume, $\hat{X}^{\mu}(\sigma)$ the embedding coordinate fields, and $\gamma_{MN}$ is the (dynamical) world-volume metric. The cosmological constant term, as usual, ensures that the world-volume metric becomes the induced metric (as algebraic equation of motion for the world-volume metric $\gamma$):
$\gamma_{MN}=\partial_{M}\hat{X}^{\mu}\partial_{N}\hat{X}^{\nu}g_{\mu\nu}.$ (1.2)
In contrast to the string, where reparameterisations and Weyl invariance can be used to gauge-fix the world-volume metric to be flat, this is not the case for the membrane. The equations of motion for the world-volume coordinate fields $\hat{X}^{\mu}(\sigma)$ are
$0$ $=\partial_{M}(\sqrt{-\gamma}\gamma^{MN}\partial_{N}\hat{X}^{\mu})+\sqrt{-\gamma}\gamma^{MN}\Gamma^{\mu}{}_{\kappa\lambda}\partial_{M}\hat{X}^{\kappa}\partial_{N}\hat{X}^{\lambda},$
${}\quad+\epsilon^{NKL}\frac{1}{3!}{F^{\mu}}_{\nu\kappa\lambda}\partial_{N}\hat{X}^{\nu}\partial_{K}\hat{X}^{\kappa}\partial_{L}\hat{X}^{\lambda},$ (1.3)
@@
Famously [17, 18], even this supposedly most simple case leads to complications, such as a continuous spectrum and spike instabilities. Quantum non-integrability of the membrane in the sense of scattering of fluctuations around a classical solution, and its relation to string theory, has recently been shown in [19]. In this letter, we will see that a large wrapping limit together with a scaling limit of the background can simplify these equations. In the special case considered here, the result will even be a classically integrable system.
@@ 1.2 The Manakov-Zakharov-Ward model
Classically integrable field theories in dimensions higher than $(1+1)$ are rare, especially those of the $\sigma$-model type, which are abundant in $(1+1)$ dimensions (see [20, 21] for recent reviews). A notable exception is the Manakov-Zakharov-Ward (MZW) model [7, 8], an integrable field theory in $(1+2)$ dimensions. It arises via dimensional reduction from the anti-self-dual Yang-Mills equations in $(2+2)$-dimensional spacetime. This reduction involves assuming that the gauge potentials are independent of one spatial coordinate, resulting in a non-Lorentz invariant system in $(1+2)$ dimensions. [^1] As a recent approach to obtain new integrable models in higher dimensions, let us mention [22].
@@
2 A decoupling limit for the membrane on T2
@@
Here, we present two (ultimately equivalent) views on a decoupling limit that can be interpreted either as a large wrapping limit around the $T^{2}$ or a non-relativistic expansion (presented in appendix A). In particular, we aim at the following, at leading order in these limits:
@@
  • membrane dynamics along $\mathcal{M}_{H}$ $\rightarrow$ MZW-model
This will involve scaling limits – both on the target space and one on the membrane world-volume.
@@
For this, let consider a general ansatz introducing scales $R$ and $\bar{R}$ for the longitudinal ($\mathcal{M}_{\tau}$, coordinates [^3] In a slight abuse of notation, we take the same indices $K,L,M,$... for the transversal coordinates $\hat{X}^{M}$ and the world-volume coordinates $\sigma^{M}$. $\hat{X}^{M}$) and transversal spaces ($\mathcal{M}_{H}$, coordinates $\hat{X}^{m}$)
$\mathrm{d}s^{2}$ $=R^{2}\eta_{MN}\mathrm{d}\hat{X}^{M}\mathrm{d}\hat{X}^{N}+\bar{R}^{2}H_{mn}\mathrm{d}\hat{X}^{m}\mathrm{d}\hat{X}^{n},$ (2.1)
$F$ $=\frac{1}{3!}R\bar{R}^{2}F_{klmN}\mathrm{d}\hat{X}^{k}\wedge\mathrm{d}\hat{X}^{l}\wedge\mathrm{d}\hat{X}^{m}\wedge\mathrm{d}\hat{X}^{N}.$ (2.2)
The scaling of $F$ has to be introduced in order to be consistent with the supergravity equations – compare (3.3).
Moreover, on the world-volume we consider arbitrary fluctuations around a static wrapped membrane in a large wrapping number limit $\omega\rightarrow\infty$:
@@
The parameter $\omega$ describes both a ’world-volume speed of light’ and the wrapping number around the torus. [^4] In principle, these parameters could be different, of course – i.e. $X^{M}=(\omega_{0}\sigma^{0},\omega_{1}\sigma^{1},\omega_{2}\sigma^{2})$. For the limit in question, we want to assume $\omega_{0}\sim\omega_{1}\sim\omega_{2}\sim\omega\rightarrow\infty$. Finite quotients could be compensated for by rescaling of the membrane coordinates $\sigma^{M}$. In the limit $\omega\rightarrow\infty$, the $\mathcal{O}(\omega^{0})$ world-volume fields $X^{\mu}(\sigma)$ are suppressed and the induced metric simplifies significantly. In particular, one can compute
$\gamma^{MN}$ $=\frac{1}{\omega^{2}R^{2}}\left(\eta^{MN}-\frac{1}{\omega}2\partial^{(M}X^{N)}+\mathcal{O}(\omega^{-2})\right),$ (2.4)
$\sqrt{-\gamma}$ $=\omega^{3}R^{3}\left(1+\frac{1}{\omega}\partial_{M}X^{M}+\mathcal{O}(\omega^{-2})\right),$
in order to compute the expansion of the equations of motion (1.3) for $X^{M}$ and $X^{M}$ at leading order
$0$ $=\omega R\left(0+\mathcal{O}\left(\omega^{-1}\right)\right)$ (2.5)
${}\quad+\frac{\bar{R}^{2}}{R}\frac{1}{3!}\eta^{KM}F_{Mklm}\partial_{N}X^{k}\partial_{L}X^{l}\partial_{R}X^{m}\epsilon^{NLR},$
[…]
${}\quad\left.+\frac{1}{2}H^{mn}F_{klnN}\partial_{K}X^{k}\partial_{L}X^{l}\epsilon^{KLN}+\mathcal{O}\left(\omega^{-1}\right)\right).$
Besides $\omega\rightarrow\infty$, the following scaling for the (dimensionless) scales $R$ and $\bar{R}$ is required, simultaneously ensuring the form (2.4) and guaranteeing that the transversal fluctations decouple from from the four-form flux at leading order:
$\frac{\bar{R}^{2}}{R^{2}\omega}\rightarrow 0.$ (2.7)
Remarkably, at leading order in the large wrapping limit the contributions from the metric fluctations (2.4) cancel completely in the equations motion (2.5).
There are several ways to interpret this limit. Let us outline two interpretations with very different consequences:
  • One can solve (2.7) with the non-Lorentzian parameterisation introduced in appendix A. For this let us introduce the dimensionless scale $c$ with
    $R^{2}=c^{2},\qquad\bar{R}^{2}=\frac{1}{c}.$ (2.8)
    In order to satisfy (2.7), the scaling of $\omega\sim c^{\alpha}\rightarrow\infty$ has to be chosen accordingly in terms of $c$. There are two choices
    • a non-relativistic limit $c\rightarrow\infty$: $\alpha>0$.
    • a hyper-relativistic (Carollian) limit $c\rightarrow 0$: $\alpha<-3$.
    Interestingly, in both cases this corresponds to large energy limits $E=\omega R\rightarrow\infty$. Although at leading order one reproduces the MZW-model, these cases will lead to additional equations of motion at the finite but subleading order. [^5] Additional contributions would arise at order $\mathcal{O}(R)<\mathcal{O}(\omega R)$ or $\mathcal{O}(\bar{R}^{2}/R)\mathcal{O}(\omega R)$.
  • One can require that all the subleading contributions are supressed, and only the leading order contributions (proportional to $\omega R$) could be potentially finite. This leads to the following additional set of constraints.
    $R\rightarrow 0,\quad\frac{\bar{R}^{2}}{R}\rightarrow 0$ (2.9)
    This is solved by a parameterisation $R\sim\bar{R}\sim\frac{1}{\omega^{\alpha}}$, with $\alpha>0$. Then this limit can be interpreted as either a large or small energy $\omega R$ limit for $\alpha>0$ or $\alpha<0$ respectively. In any case, this is a small volume regime in target space – both scales $R,\bar{R}\rightarrow 0$.
@@ 3.1 A toy model background
In order to obtain the MZW-model from the transverse equations of motion at leading order, like (2.6), we specify $\mathcal{M}_{H}$ to be a Lie group $G$.
$\mathrm{d}s^{2}$ $=R^{2}\eta_{MN}\mathrm{d}\hat{X}^{M}\mathrm{d}\hat{X}^{N}+\bar{R}^{2}\kappa_{ab}e^{a}e^{b}$ (3.1)
$F$ $=-R\bar{R}^{2}\frac{1}{3!}f_{abc}e^{a}\wedge e^{b}\wedge e^{c}\wedge v.$
$e^{a}={e_{m}}^{a}\mathrm{d}\hat{X}^{m}$ are the Maurer-Cartan one-forms on $G$, $f_{abc}$ are the structure constants for the Lie algebra $\mathfrak{g}$ of $G$, and $\kappa$ is the Killing form on $\mathfrak{g}$. $v$ is a one form on $\mathcal{M}_{\tau}$. We choose $v=\mathrm{d}\hat{X}^{0}$ or $v=\mathrm{d}\hat{X}^{i}$, for $i=1,2$. This choice will be related to the choice of $v$ in the MZW-model (1.4). In the same way as in the previous section, at leading order the membrane equations of motion, now expressed in terms of currents $j^{a}={e_{m}}^{a}\mathrm{d}X^{m}$, are
$0$ $=\partial^{N}j_{N}^{a}-\frac{1}{2}{f^{a}}_{bc}j^{b}_{M}j^{c}_{N}v_{L}\epsilon^{MNL},$ (3.2)
for the expansion (2.3). We recognise the MZW-model’s equation of motion (1.4) for the transverse directions. The longitudinal equations of motion completely decouples at leading order as in (2.5).